Presented below is information related to BLANK radios for the BLANK Company for the month of July.

Date

Transaction

Units In

Unit Cost

Total

Units Sold

Selling Price

Total

July 1 Balance 130 $3.90 $  507
6 Purchase 1,040 4.00 4,160
7 Sale 390 $6.70 $ 2,613
10 Sale 390 7.00 2,730
12 Purchase 520 4.30 2,236
15 Sale 260 7.10 1,846
18 Purchase 390 4.40 1,716
22 Sale 520 7.10 3,692
25 Purchase 650 4.38 2,847
30 Sale 260 7.20 1,872
   Totals 2,730 $11,466 1,820 $12,753

Assuming that the periodic inventory method is used, compute the inventory cost at July 31 under each of the following cost flow assumptions. (Round answers to 0 decimal places, e.g. 6,578.)

(1) FIFO.
(2) LIFO.
(3) Weighted-average.

(1)
FIFO

(2)
LIFO

(3)
Weighted-Average

Ending Inventory at July 31 $

Entry field with incorrect answer now contains modified data

$

Entry field with incorrect answer now contains modified data

$

Entry field with incorrect answer now contains modified data
Answer #1

Closing inventory units = 2,730 - 1,820 = 910.

FIFO Method

>> Closing inventory = ( 650 * $ 4.38 ) + ( 260 * $ 4.40 ) = $ 3,991.

LIFO Method

>> Closing Inventory = ( 130 * $ 3.90 ) + ( 780 * $ 4 ) = $ 3,627.

Weighted average method.

Weighted average cost = $ 11,466 / 2,730 = $ 4.2

>> Closing Inventory = 910 * $ 4.2 = $ 3,822.